OpenStudy (anonymous):

Find the product, if possible. AB, if A= [0 -5 10] B= [-9] [-1 6 -9] [0] , [4]

OpenStudy (anonymous):

0*-9+-5*0+10*4=40 -1*-9+6*0+-9*-4=45 so AB is 40 45